OpenStudy (suwhitney):

Accounting Question: Compute annual depreciation for the first and second years under the declining-balance method...

OpenStudy (suwhitney):

Weller Company acquires a delivery truck at a cost of $69,300. The truck is expected to have a salvage value of $2,300 at the end of its 5-year useful life. Assuming the declining-balance depreciation rate is double the straight-line rate. Compute annual depreciation for the first and second years under the declining-balance method.

OpenStudy (suwhitney):

I cannot seem to get this question right, any help would be MUCH appreciated!

OpenStudy (suwhitney):

@johnweldon1993 are you as good at accounting as you are math? :P

OpenStudy (johnweldon1993):

Well from what I remember from accounting Your straight-line rate would be 1/5 = 20% since it has a 5 year useful life so double that is 40% so lets find out the depreciation after 1 year \[\large 69300 \times .40 = ?\]

OpenStudy (johnweldon1993):

Still there @suwhitney ?

OpenStudy (suwhitney):

So sorry @johnweldon1993! I got pulled away for a second.. 69300x.40=27720 :) And thats for year 1? Correct?

OpenStudy (suwhitney):

omg i figured the rest out! :)

OpenStudy (johnweldon1993):

No problem :) and yes...that is the annual depreciation for year 1 Now we need the book value after year 1 which is just the original - the depreciation... 69300 - 27720 = ? From that answer, we can then figure out the depreciation after year 2 by multiplying that answer by .40

OpenStudy (suwhitney):

16632!

OpenStudy (suwhitney):

Yayyy!! Thank you so much! You're a genius :P

OpenStudy (johnweldon1993):

Oh nvm :D haha awesome ^_^ :P